OpenStudy (solomonzelman):

A little tutoring try not to interrupt. Watch, but don't type. laurenlauren12345

OpenStudy (solomonzelman):

@laurenlauren12345

OpenStudy (anonymous):

I am here

OpenStudy (solomonzelman):

Ok, tell me what you know and what you don;t know so that I know what to tach you. Do you know how to graph a line that looks just like y=3x+5 ?

OpenStudy (anonymous):

I know y=mx+b ! Oh and I am only in 8th grade by the way

OpenStudy (solomonzelman):

Ok, show me then how would you graph a line, y=3x+5

OpenStudy (anonymous):

Ummmm....... U put 3x on the x and the 5 on the y

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

YOu choose any random x-values and find what would y-value be with a certain x-value.

OpenStudy (solomonzelman):

For example (I am choosing any random number, lets say) -2 y=3x+5 plug in -2 for x y=3(-2)+5 y=-6+5 y=1 so you know that when x=-2 y=-1 correct?

OpenStudy (anonymous):

Sorry but I am a little confused

OpenStudy (solomonzelman):

Okay, see how I am pluging in -2 for x?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

and see how I got y=-1 when I plugged -2 for x?

OpenStudy (anonymous):

I am kinda confused on that one

OpenStudy (solomonzelman):

The equation is y=3x+5 right?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

We plugged in -2 instead of x, like this y=3x+5 y=3(-2)+5 okay?